Model 2 – The Alkali Meta Atomic Number 3 Core Charge Lithium Sodium I Potassium 119 9. Describe the relationship of 1: elati what is the core charge of these elements?
Answer #1

Core charge = no of protons - Inner shell e Lithium = 3-2 = 1 11- 10 = +1 Badium - potanium - 19-18 = 71 = [1 3²/23 wa = 1152